OpenStudy (anonymous):

what is the vertex of (-4x^2)-x+5

OpenStudy (anonymous):

Do you mean Roots??????

OpenStudy (anonymous):

nope vertex.. i got -1/8 for x i just need why now

OpenStudy (anonymous):

Use -b/2a.

OpenStudy (anonymous):

i did thats how i got -1/8

OpenStudy (anonymous):

That would be your x value. Now plug that back in to get your y value.

OpenStudy (anonymous):

So you want to know why we apply -b/2a and what is its significance?????

OpenStudy (anonymous):

Okay i plugged it back in and i ended up getting 5.062 but i think it wants it in fraction form

OpenStudy (anonymous):

Fraction form would be 81/16

OpenStudy (anonymous):

*it refers to Wiley Plus

OpenStudy (anonymous):

\[(-4x^2)-x+5\] \[-4(-\frac{1}{8})^2-(-\frac{1}{8})+5\]

OpenStudy (anonymous):

ok great thanks zordoloom

OpenStudy (anonymous):

yep.

OpenStudy (anonymous):

\[-\frac{1}{16}+\frac{1}{8}+5\] \[\frac{1}{16}+5\] \[5\tfrac{1}{16}\]or if you prefer \[\frac{81}{16}\]